Exercise 2.54

Answers

(a)
pk = ppk−1 + qpk+1 with boundary condition p0 = 1.
(b)
Let Aj be the event that the drunk reaches k before reaching − j. Then, Aj ⊆ Aj+1 since to reach − (j + 1) the drunk needs to pass − j. Note that ⋃ ⁡ j=1∞Aj is equivalent to the event that the drunk ever reaches k, since the complement of this event, namely the event that the drunk reaches − j before reaching k for all j implies that the drunk never has the time to reach k.

By assumption, P(⋃ ⁡ j=1∞Aj) = lim ⁡ n→+∞P(An). P(An) can be found as a result of a gambler’s ruin problem.

If p = 1 2,

P(An) = n n + k → 1.

If p > 1 2,

P(An) = 1 − (q p ) n 1 − (q p ) n+k → 1.

If p < 1 2,

P(An) = 1 − (q p ) n 1 − (q p ) n+k → (p q )k.

User profile picture
2021-12-05 00:00
Comments